OpenStudy (anonymous):

Will medal and fan. A paint sprayer covers 2 1/2 ft of wall in 1/8 min. Complete the unit rate. How many feet per minute?

OpenStudy (anonymous):

20 feet per minute

OpenStudy (anonymous):

Thank you

OpenStudy (texaschic101):

proportion time 2 1/2 = 5/2 (5/2) / (1/8) = x/ 1 --- 5/2 ft to 1/8 min = x ft to 1 minute) cross multiply 1/8x = 5/2 x = (5/2) / (1/8) x = 5/2 * 8/1 x = 40/2 x = 20 minutes

OpenStudy (anonymous):

Thank you very much, that makes a lot more sense now